OpenStudy (giitter):

Help me I give medal and fan if you answer correctly :D Enjoy answering my questions XD

OpenStudy (anonymous):

Ok this is the first one XD Which expression could be used to solve this problem? John reads the newspaper for 10 minutes every morning. He reads a book for 15 minutes every evening. How much time does John spend reading in 5 days? A. 5 × (10 + 15) B. 5 × (15 – 10) C. 15 × (10 + 5) D. 15 × (10 – 5)

OpenStudy (anonymous):

Which strategy could be used to solve 14 × 32? A. (14 × 3) + (14 × 2) =? B. (14 + 3) + (14 + 2) =? C.(4 × 32) + (10 × 32) =? D. (4 + 32) + (10 + 32) =?

OpenStudy (xmissalycatx):

Okay, for the first question, what do you think it is?

OpenStudy (anonymous):

I dunno

OpenStudy (anonymous):

\[5\times(10+15)\] or A

OpenStudy (xmissalycatx):

@JavaScriptCoder Do NOT give direct answers! It is against the code of conduct!

OpenStudy (anonymous):

I know that why is it that answer @JavaScriptCoder? I will not trust you unless you give me the reason WHY its the answer

OpenStudy (xmissalycatx):

It is the correct answer because if you have 10 + 15, you're just adding the amount of hours/minutes/seconds etc.. of which you are reading together. The reason you multiply them by 5 is because it is 5 total days.

OpenStudy (anonymous):

Ok thanks and I fanned you already

OpenStudy (xmissalycatx):

Yeah I know. :P

OpenStudy (anonymous):

XD

OpenStudy (xmissalycatx):

OH SNAP GURL DIDN'T SEE THERE WAS A 2ND QUESTION XD

OpenStudy (anonymous):

XD

OpenStudy (xmissalycatx):

I sent you a message. I have to go. :P

OpenStudy (anonymous):

bye